OpenStudy (anonymous):

Use implicit differentiation find dy/dx of xy^1/3+y=10

OpenStudy (akashdeepdeb):

\[xy^{\frac{1}{3}} + y = 10\] Differentiate both sides. \[\frac{d}{dx} (xy^{\frac{1}{3}} + y ) = \frac{d}{dx} (10)\] \[\frac{d}{dx} (xy^{\frac{1}{3}} ) + \frac{d}{dx} (y) = 0\] Apply product rule and find \(\frac{dy}{dx}\). Getting this? :)

OpenStudy (anonymous):

U can simply do the derivative of xy^1/3 using chain rule. For y you would just write dy/dx. Then 10 is a constant so it would just be 10. At last, you solve for dy/dx.

OpenStudy (anonymous):

ya Im getting lost somewhere after I've applied the product rule. is this right? \[(1)(y ^{1/3}) + (x)(1/3y ^{-1/6})(dy/dx) +dy/dx=0\]